OpenStudy (anonymous):

Solve x2 = 14x – 49.

OpenStudy (anonymous):

x^2 = 14x - 49

OpenStudy (anonymous):

@KCB it uses the same concepts as the previous one so try it urself if u don't get it we will help u

OpenStudy (anonymous):

I think it would be x^2 +14x +49 =0 for the first step

OpenStudy (anonymous):

Yes ! and we have a Quadratic equation:) Do you know that ?:)

OpenStudy (anonymous):

Yes @KCB ur correct

OpenStudy (anonymous):

I do now lol. so how would i find the next step ? do i have to make an equation with numbers suming up to 49 or ??

OpenStudy (anonymous):

x^2+7x+7x+49=0 x(x+7)+7(x+7)=0 (x+7)(x+7)=0 x=-7

OpenStudy (anonymous):

okay thanks that doesnt seem to hard. Im going to post one more and solve it . WIll you let me know if im correct @UditKulka

OpenStudy (anonymous):

@UditKulka :Are you sure ???:)!!!! I think you re falllll !!!:(

OpenStudy (anonymous):

@E.ali how would u solve it ?

OpenStudy (anonymous):

We should use from the Quadratic equation formula ! We have 2 answers : X1 and X2 :) K ?:)

OpenStudy (anonymous):

x= -7 or x= +7

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i made a mistake i am extremely sorry for that

OpenStudy (anonymous):

the equation is x^2-14x+49

OpenStudy (anonymous):

@UditKulka : Thank you friend !

OpenStudy (anonymous):

x2-7x-7x+49 x(x-7)-7(x-7) (x-7)(x-7) x=+7-----------------So only only one sollution of x in this equation i suppose @E.ali

OpenStudy (anonymous):

so besides leaving out the other answer @UditKulka is correct ? @E.ali

OpenStudy (anonymous):

Wait friends ! Step by step ! Now look at me !:

OpenStudy (anonymous):

the equation is Solve x^2 = 14x – 49. idk im confuse d

OpenStudy (anonymous):

@KCB There are different methods of solving Quadratics and first of all don't panic